What mass of solute is contained in 20.0 ml of a 2.09 M barium acetate solution?
1
Expert's answer
2019-06-05T03:58:23-0400

Ci=n/V --> n=Ci*V

m=n*M --> m=Ci*V*M

M(BaC4H6O4)=137+12*4+1*6+16*4=255g/mol

m(BaC4H6O4)= 2.09*0.02*255=10.66g


Answer: 10.66 g